# The Role of Inclusion, Control, and Ownership in Workplace AI-Mediated Communication

### v1

@@ Abstract.
Large language models (LLMs) can exhibit social biases. Given LLMs’ increasing integration into workplace software, these biases may impact workers’ well-being and may disproportionately impact minoritized communities. This short paper investigates how co-writing with an LLM impacts three measures related to user’s well-being: feelings of inclusion, control, and ownership over their work. In an online experiment, participants wrote hypothetical job promotion requests to their boss and using either hesitant or self-assured auto-complete suggestions from an LLM. Afterward, participants reported their feelings of inclusion, control, and ownership. We found that the style of the AI model did not impact perceived inclusion. Furthermore, individuals with higher perceived inclusion also perceived greater agency and ownership, an effect more strongly impacting participants of minoritized genders. Lastly, feelings of inclusion can mitigate a loss of control and agency when accepting more AI suggestions. Future work should explore feelings of inclusion in AI-written communication.
@@ 1. Introduction
Large language models (LLMs) have the potential to transform workplace communication. These models are already seeing widespread use assisting in common workplace tasks like writing emails. Recent scholarship has shown that workers can use LLMs at work to increase their productivity and produce higher-quality work (Noy and Zhang, 2023). However, as others have noted, one potential challenge is that language models can reflect social biases in the language they produce (Weidinger et al., 2021; Bender et al., 2021). While the existence of such biases is well-documented, we know far less about how these biases affect individual users. In particular, when AI tools—which may reflect prevailing social norms—are used in the workplace, they may inflict a “psychological tax” for individuals who do not align with these norms (Weidinger et al., 2021), affecting their perceived inclusion, control, and ownership over their work output.
Feelings of inclusion, control, and ownership are core to the human experience (Ryan and Deci, 2017). People who feel included develop a sense of belonging (Sargent et al., 2002) that can significantly improve their psychological health (Lambert et al., 2013; Reker and Wong, 2013). For example, people with a strong sense of belonging have shown increased resilience (Sargent et al., 2002), and a strong sense of belonging in the workplace can mediate negative feelings like anxiety and depression (Hartley, 2011; Cockshaw and Shochet, 2010). Furthermore, employees with a strong sense of belonging are more motivated to be productive and reach their fullest potential (Isham et al., 2020; Walton et al., 2012; Waller, 2020). Similarly, perceptions of control and ownership positively contribute to employees’ workplace satisfaction and job performance (Dyne and Pierce, 2004; Eatough and Spector, 2014; Pierce et al., 2004).
Given the relationship between inclusion, control, and ownership and their implications for individual well-being in professional environments, it is crucial to explore how elements of workplace AI-mediated communication (AI-MC) (Hancock et al., 2020) can impact these feelings, especially for users’ from minoritized groups. In particular, we focus here on the potential impact of AI writing assistants that provide auto-complete suggestions as the user is typing text, for example while writing an email. In this context, we ask the following questions:
@@
In the current study, we investigate the impact of a specific communication style— assertiveness—on minoritized gender groups. We build on a rich history examining gender disparities, particularly in the workplace where assertive language can be very consequential (Amanatullah and Morris, 2010; Smith, 2002) and associated with better work outcomes (Mazei et al., 2015; Amanatullah and Morris, 2010). While there are many cultural differences that guide expectations of gender presentation in the workplace, scholars in gender-based communication describe male talk in Western workplaces as being assertive, powerful, and authoritative (Mullany, [n. d.]; Merchant, 2012; Thimm et al., 2008). If writing assistants making suggestions in a more assertive style are not as well-aligned with the communication style of women and members of other minoritized gender groups, the discrepancy in style alignment could cause those individuals to feel a lower sense of inclusion, control, and ownership.
Specifically, this short paper addresses our research questions using an online experiment where participants completed a hypothetical, but common and consequential workplace writing task, namely writing to their manager asking for a promotion. Participants received auto-complete suggestions from one of two LLM-powered AI writing assistants that offered suggestions in different styles: assertive or hesitant. We then asked participants about their feelings of inclusion, control, and ownership in the writing experience.
We found that the intervention—getting suggestions in assertive or hesitant style from an AI assistant—did not impact participants’ perceptions of inclusion, though it did impact their perceptions of control and ownership. Surprisingly, we did not find the assistant style treatments directly resulted in differences between genders in our main measures. We do find evidence that gender plays some role: namely that minoritized genders’ perception of control and ownership are more greatly impacted by their (likely pre-existing) feelings of inclusion than men’s. In light of these findings, we offer a new conceptual model for understanding the impact of these factors on people’s feelings of overall agency in AI-mediated environments.
@@ 2. Background
While AI-powered intelligent writing assistants are a relatively recent development, there is already a rich set of research contributions on our interactions with writing assistants and the biases they may introduce.
As large language models became more robust, with the ability to inspire ideas, revise text, and generate short stories, the field has shifted toward viewing other dimensions of the writing process (Singh et al., 2022; Bhat et al., 2023). There is a growing interest in viewing writing assistants powered by large language models as co-authors and studying the interaction between users and these systems (Lee et al., 2022; Mirowski et al., 2023). Recent scholarship in this area has focused on understanding how writers evaluate and integrate the suggestions provided by these language models into their cognitive writing processes (Bhat et al., 2023). Writers proactively engage with system-generated content, making deliberate efforts to integrate it into their writing (Singh et al., 2022). However, there is evidence that perceptions of ownership can differ. Inexperienced writers perceive less ownership than experienced writers during the co-writing process (Mieczkowski, 2022). Here, we expand the understanding of how these systems can best support and enhance the writing process by demonstrating the importance of inclusion when interacting with an AI writing assistant.
As large language models like GPT (Vaswani et al., 2023; Bommasani et al., 2022) are increasingly integrated into our society, there is growing concern among AI ethicists regarding their potential harms, including in the context of auto-complete suggestions. Concerns have been raised regarding the risks of misinformation, privacy violations, socioeconomic harms, and, most salient to our work here, representational harms (Weidinger et al., 2021). Large language models are trained on large corpuses of data which often reflect historic and systemic injustices. For example, a language model is more likely to produce the phrase “he is a doctor” than “she is a doctor”, assuming the subject’s gender or ethnic identity (Weidinger et al., 2021). As a result, these models may produce demeaning language and stereotypes which can be especially harmful for some with intersecting identities. Specifically in our context, biases in auto-complete suggestions have been shown to shift the use of language and influence content (Hohenstein et al., 2023; Poddar et al., 2023; Jakesch et al., 2023). AI suggestions even shift people’s opinions (Jakesch et al., 2023). Extending this body of work, here we show how AI biases may impact the perceptions of people using these technologies.
@@ 3.1. Writing App
We used a custom experimental platform combining a rich-text editor and a writing assistant, as seen in Figure 1. The platform code was made available by researchers (Jakesch et al., 2023). The suggestions work such that when the participant pauses typing, the system displays suggestions that are 20 words or less. The user can then choose to accept the suggestion by pressing tab or right arrow key, or they can ignore the suggestion by continuing to type. Pressing tab accepts the subsequent word in the suggestion, and a user can press multiple times to accept the full suggestion.
@@ 3.2. Participant Recruitment
We recruited 750 participants from a gender-balanced sample of English-speaking US-based adults on Prolific. The sample size was calculated using a two-way ANOVA with a power of 80% based on a pre-test survey. Tp ensure response quality, we required participants to have approval ratings greater than 95% on Prolific. We excluded participants who did not state their gender, and manually verified that each writing sample was faithful to the writing task. In the end, we had usable data from 738 participants.
Our experiment placed emphasis on two distinct gender categories: individuals with significant societal influence, predominantly men, and minoritized genders who typically encounter societal disparities. While describing our results we will refer to Women, Non-Binary or third gender, and “preferred to describe” (also known as Self-Described) as WNBSD participants. Of the 738 participants whose data is used, 48.9% identified as men, and 51.1% as WNBSD: 47.7% identified as women, 3% identified as non-binary or third gender, and 0.4% preferred to describe their gender.
The majority of the participants were between 25-34 years old (56%), identified as white (61.9%), received a bachelor’s degree (38.2%), reported working a full-time job (59.4%), and reported working entirely in person (40.6%).
@@ 3.3. Measures
AI reliance. Prior work on how users interact with AI writing assistants uses the measure of “mutuality” to describe the level of interaction the writer has with the AI assistant (Mieczkowski, 2022; Lee et al., 2022). The measure uses a score from 0 to 1, where 1 means the assistant wrote the entire text (reflecting complete reliance on the assistant) and 0 means the human wrote the entire text (no reliance). We use the term AI reliance for this measure, instead of mutuality, as we believe this concept better captures the construct used.
Inclusion. We draw on the concept of self-extension, which explains how objects become integral to one’s self-concept (Belk, 2013), to develop the notion of inclusion in AI-mediated communication. We focus on identity self-extension where the technology is a reflection of the user (Ross and Bayer, 2021; Mieczkowski, 2022). This literature suggests that if a user feels the technology reflects themselves, they will feel included. We therefore measure inclusion with the following statements: the writing assistant was made for people like me and the writing assistant sounded like I would write myself.
Control. Prior work has shown that feeling in control is not limited to one action or outcome. Feeling in control can encompass engaging in an action, control over the outcome, or both (Singh et al., 2022; Mieczkowski, 2022). For example, some users may a feel lack of control since they cannot choose which suggestions they are seeing (control during the writing process). However, the fact that they can select which suggestions are incorporated into the message can impact their sense of control over the process and the final version (Bhat et al., 2023). We therefore measure by asking participants how much control did you feel over the process of writing the message and how much control did you feel over the final version of the message (Mieczkowski, 2022).
Ownership. One aspect of ownership involves ”mineness”– the idea that a specific target belongs to an individual. The target can include physical objects but can also encompass intangible objects such as words and thoughts. In our experiment, we consider two targets– the message itself and the message style. Therefore, we also investigate ownership over the writing content and style. Our questions included to what extent do you feel like the message you wrote is yours and thinking back on the message writing activity, how much did the message sound like you (Mieczkowski, 2022).
@@ 4. Results
Our results expose the impact of writing style on perceptions of inclusion, control, and ownership and agency and hint at potential differences between genders in that respect. Our primary, preregistered analysis finds some significant effects of AI writing style on perceptions of control and ownership, but not on inclusion. A follow-up, exploratory analysis suggests that as AI reliance increases, perceptions of inclusion mediate the loss of agency. Additionally, the analysis suggests that heightened perceptions of inclusion lead to increased agency, particularly among minoritized genders.
We first provide an overview of the effects of using the AI auto-complete suggestions on the writing task. We examined three components of the writing task: the length of text written by participants, their writing time, and their degree of AI reliance. We performed a two-way ANOVA to compare the means between the three treatment groups (self-assured AI writing style, hesitant AI writing style, and control) and two gender groups (men and WNBSD participants). Overall, participants who wrote with the assistant wrote longer emails (M=160.18, SD=82.16) and spent less time on the task (M=303.82, SD=218.53) than participants in the control group (email length M=95.53, SD=35.64; time on task M=363.66, SD=292.68). The differences are statistically significant (length: ANOVA F=67.86, p¡0.00001; time: ANOVA F=6.35, p¡0.002). Post-hoc tests did not observe any differences between the two AI conditions in email length and time on task. In contrast, we did observe significant impacts of AI writing style on AI reliance: participants who wrote with the self-assured writing style model relied on the model more (M=0.55, SD=0.34) than participants who wrote with the hesitant writing style model (M=0.45, SD=0.35). These differences were also significant (t=3.16, p¡0.002). We did not observe gender effects or interaction effects between gender and the AI treatment on any of these metrics.
@@ 4.1. Inclusion
We examine the impact of AI writing style on our key measures, beginning with perceptions of inclusion. Figure 2 shows, for each gender condition (different bars) the frequency of different survey responses for the inclusion statements: the writing assistant was made for people like me (“Assistant Was Made for You” on the left), and the writing assistant sounded like I would write myself (“Assistant Sounds Like You” on the right). Figure 2a (top-left) shows that in the condition with the self-assured AI writing style, 68.3% of the men and 68.5% of the WNBSD participants either agreed or strongly agreed (the blue and dark blue shades on the right of the bars) that the writing assistant was made for people like them. As the figure also shows, in the hesitant AI writing style group, (the bottom-left Figure 2c), 65.1% of men and 62.9% of WNBSD participants either agreed or strongly agreed with this statement.
The responses to the other statement, whether the assistant aligned with the participants’ writing style, are shown on the right side of Figure 2. In general, participants more readily agreed with this statement compared to the earlier “made for you” statement. Figure 2b (top-right) shows the participants who wrote with the self-assured writing style model; of the participants who received this treatment, 50% of the men and 51.7% of the WNBSD participants either agreed or strongly agreed that the assistant sounded like how they would write themselves. In the condition with the hesitant AI writing style, (bottom-right Figure 2d), 42.9% of the men and 45.7% of the WNBSD participants either agreed or strongly agreed. A two-way ANOVA analysis revealed no significant effects for gender, AI treatment group, or the interaction between gender and group across all inclusion questions.

### v2

@@ Abstract.
Given large language models’ (LLMs) increasing integration into workplace software, it is important to examine how potential biases they introduce can impact workers. Stylistic biases in the language suggested by LLMs may cause feelings of alienation and result in increased labor for individuals or groups whose style does not match. We examine how such writer-style bias impacts inclusion, control, and ownership over the work when co-writing with LLMs. In an online experiment, participants wrote hypothetical job promotion requests using either hesitant or self-assured auto-complete suggestions from an LLM and reported their subsequent perceptions of inclusion, control, and ownership. We found that the style of the AI model did not impact perceived inclusion. However, individuals with higher perceived inclusion did perceive greater agency and ownership, an effect more strongly impacting participants of minoritized genders. Feelings of inclusion mitigated a loss of control and agency when accepting more AI suggestions.
@@ 1. Introduction
Large language models (LLMs) have the potential to transform workplace communication but can also exacerbate existing societal biases. LLMs are already seeing widespread use in assisting in common workplace tasks like writing emails and authoring documents. Recent scholarship has shown that workers can use LLMs to increase their productivity and produce higher-quality work (Noy and Zhang, 2023). Despite these potential benefits, LLMs also have drawbacks, including the potential to reproduce social biases (Weidinger et al., 2021; Bender et al., 2021). For instance, when LLMs generate text about particular identity groups, they may display social biases in the content and style of the resulting text (Wan et al., 2023), causing representational harms through stereotyping and demeaning language (Weidinger et al., 2021; Shelby et al., 2023). In addition to representational harms, another category of harm and the one we focus on in this work, is quality-of-service harm (Shelby et al., 2023; Bird et al., 2020), in which systems unequally serve different groups, and which may “result in feelings of alienation, increased labor, and service or benefit loss” for people of different identities (Shelby et al., 2023). For example, if an LLM produces text that is stylistically better aligned with one gender group over another, the result is a model that disparately benefits certain users over others.
Such stylistic biases, whether real or perceived by users, can also impact how individuals and groups of end-users feel about and interact with AI tools. Following the previous example, when an LLM’s text style better suits one group over another, this may impact users’ feelings of inclusion when using that system. Inclusion could affect other psychological factors like users’ sense of the human experiences of control and ownership (Ryan and Deci, 2017) during the co-writing process. For example, lower perceptions of inclusion may lead users to accept fewer AI suggestions; prior work has shown the relationship between accepting AI suggestions and feelings of agency and control (Mieczkowski, 2022).
Given the potential importance of feelings of inclusion on people’s use of AI tools, we set out to examine the relationship between writer-style bias and perceptions of inclusion, control, and ownership while co-writing with AI. In particular, we focus here on the potential impact of AI writing assistants that provide auto-complete suggestions as the user is typing text, for example, while writing an email. In this context, we ask the following questions:
@@
More specifically, in this work, we investigate the impact of a specific communication style—assertiveness—on writers and its impact on minoritized gender groups. We build on a rich history examining gender disparities, particularly in the workplace where assertive language can be very consequential (Amanatullah and Morris, 2010; Smith, 2002) and associated with better work outcomes (Mazei et al., 2015; Amanatullah and Morris, 2010). While there are many cultural differences that guide expectations of gender presentation in the workplace, scholars in gender-based communication describe male talk in Western workplaces as being assertive, powerful, and authoritative (Mullany, [n. d.]; Merchant, 2012; Thimm et al., 2008). If writing assistants making suggestions in a more assertive style are not as well-aligned with the communication style of women and members of other minoritized gender groups, the discrepancy in style alignment could cause an additional burden for women and gender minorities by increasing writing effort and task completion time.
This short paper addresses the research questions above using an online experiment where participants completed a hypothetical but common and consequential workplace writing task:writing to a manager to ask for a promotion. Participants received auto-complete suggestions from one of two LLM-powered AI writing assistants that offered suggestions in different styles: assertive or hesitant. We then asked participants about their feelings of inclusion, control, and ownership in the writing experience.
We found that the intervention—getting suggestions in an assertive or hesitant style from an AI assistant—did not impact participants’ perceptions of inclusion. However, it did impact their perceptions of control and ownership. We did not find the assistant style treatments directly resulted in gender differences in our main measures. We do find evidence that gender plays some role: minoritized genders’ perceptions of control and ownership are more greatly impacted by their (likely pre-existing) feelings of inclusion than men’s. In light of these findings, we offer a new conceptual model for understanding the impact of task- and user-style alignment on people’s feelings of overall agency in AI-mediated environments.
@@ 2. Background
As LLMs become more robust with the ability to inspire ideas, revise text, and generate short stories, the writing assistants powered by large language models are increasingly seen as co-writers. There is a growing body of work studying the interaction between users and these co-writing systems (Lee et al., 2022; Mirowski et al., 2023). Recent scholarship in this area has focused on understanding how writers evaluate and integrate the suggestions provided by LLMs into their cognitive writing processes (Bhat et al., 2023), and how writers proactively engage with system-generated content (Singh et al., 2022).
Along with their potential benefits, however, these models carry potential risks — risks that may not be equally distributed across different user groups. Prior work has categorized harms that include social systems harms, allocative harms, representational harms, quality of service harms, and interpersonal harms (Weidinger et al., 2021; Shelby et al., 2023; Weidinger et al., 2022). In the context of auto-complete suggestions, prior work has shown that societal biases can cause information harms (a subset of social systems harms) by influencing what content is written (Poddar et al., 2023; Jakesch et al., 2023), and even shifting writers’s attitudes, potentially without their awareness (Jakesch et al., 2023; Kidd and Birhane, 2023).
Quality-of-service harms are another concern identified by researchers. These harms reflect developer’s choices and result in unintended performance disparities based on identity (Shelby et al., 2023; Bird et al., 2020). For example, prior work has shown that audio speech recognition (ASR) systems have significantly higher error rates for Black speakers than White speakers (Koenecke et al., 2020). These error rates can have psychological and behavioral impacts on Black users, such as feelings of frustration and disappointment, and can cause users to have to modify their linguistic patterns (Mengesha et al., 2021). Such quality-of-service harms can lead to alienation and result in greater labor burdens and loss of benefits for marginalized communities (Shelby et al., 2023). In this short paper, we extend research on this topic to consider writer-style biases in auto-complete suggestions and their potential to cause differential impacts on inclusion, control, and ownership for different social groups, with a specific focus on gender.
@@ 3.1. Writing App
We used a custom experimental platform combining a rich-text editor and a writing assistant, as seen in Figure 1. was made available by researchers (Jakesch et al., 2023). The suggestions work such that when the participant pauses typing, the system displays suggestions that are 20 words or less. The user can then choose to accept the suggestion by pressing tab or right arrow key, or they can ignore the suggestion by continuing to type. Pressing tab accepts the subsequent word in the suggestion, and a user can press multiple times to accept the full suggestion.
@@ 3.2. Participant Recruitment
We recruited 750 participants from a gender-balanced sample of English-speaking US-based adults on Prolific. The sample size was calculated using a two-way ANOVA with a power of 80% based on a pre-test survey. To ensure response quality, we required participants to have approval ratings greater than 95% on Prolific. We excluded participants who did not state their gender and manually verified that each writing sample was faithful to the writing task. In the end, we had usable data from 738 participants.
Our experiment emphasized two distinct gender categories: individuals with significant societal influence, predominantly men, and minoritized genders who typically encounter societal disparities. While describing our results, we will refer to Women, Non-Binary or third gender, and “preferred to describe” (also known as Self-Described) as WNBSD participants. Of the 738 participants whose data is used, 48.9% identified as men, and 51.1% as WNBSD: 47.7% identified as women, 3% identified as non-binary or third gender, and 0.4% preferred to describe their gender. The majority of the participants were between 25-34 years old (56%), identified as white (61.9%), received a bachelor’s degree (38.2%), reported working a full-time job (59.4%), and reported working entirely in person (40.6%).
@@ 3.3. Measures
AI reliance. Prior work on how users interact with AI writing assistants uses the measure of mutuality to describe the level of interaction the writer has with the AI assistant (Mieczkowski, 2022; Lee et al., 2022). However, in this work, we are interested in the final text output, not the event-based interactions. We introduce a new measure, AI reliance, as we believe this better intuitively captures the construct. AI reliance is the fraction of AI-written characters in the entire text. The measure uses a score from 0 to 1, where 1 means the assistant wrote the entire text (reflecting complete reliance on the assistant) and 0 means the human wrote the entire text (no reliance).
Inclusion. We draw on the concept of self-extension, which explains how objects become integral to one’s self-concept (Belk, 2013), to develop the notion of inclusion in AI-mediated communication. We focus on identity self-extension where the technology is a reflection of the user (Ross and Bayer, 2021; Mieczkowski, 2022). This literature suggests that if a user feels the technology reflects themselves, they will feel included. We therefore measure inclusion with the following statements: the writing assistant was made for people like me, and the writing assistant sounded like I would write myself.
Control. Prior work has shown that feeling in control is not limited to one action or outcome. Feeling in control can encompass engaging in an action, control over the outcome, or both (Singh et al., 2022; Mieczkowski, 2022). For example, some users may feel a lack of control since they cannot choose which suggestions they are seeing (control during the writing process). However, the fact that they can select which suggestions are incorporated into the message can impact their sense of control over the process and the final version (Bhat et al., 2023). We therefore measure by asking participants how much control did you feel over the process of writing the message and how much control did you feel over the final version of the message (Mieczkowski, 2022).
Ownership. One aspect of ownership involves "mineness"– the idea that a specific target belongs to an individual. The target can include physical objects but can also encompass intangible objects such as words and thoughts. In our experiment, we consider two targets– the message itself and the message style. Therefore, we also investigate ownership over the writing content and style. Our questions included to what extent do you feel like the message you wrote is yours and thinking back on the message writing activity, how much did the message sound like you (Mieczkowski, 2022).
@@ 4. Results
Our results expose the impact of writing style on perceptions of inclusion, control, ownership, and agency and hint at potential differences between genders in that respect. Our primary, preregistered analysis finds some significant effects of AI writing style on perceptions of control and ownership but not on inclusion. A follow-up, exploratory analysis suggests that perceptions of inclusion mediate the loss of agency as AI reliance increases. Additionally, the analysis suggests that heightened perceptions of inclusion lead to increased agency, particularly among minoritized genders.
We first provide an overview of the effects of using the AI auto-complete suggestions on the writing task. We examined three components of the writing task: the length of text written by participants, their writing time, and their degree of AI reliance. We performed a two-way ANOVA to compare the means between the three treatment groups (self-assured AI writing style, hesitant AI writing style, and control) and two gender groups (men and WNBSD participants). Overall, participants who wrote with the assistant wrote longer emails (M=160.18, SD=82.16) and spent less time on the task (M=303.82, SD=218.53) than participants in the control group (email length M=95.53, SD=35.64; time on task M=363.66, SD=292.68). The differences are statistically significant (length: ANOVA F=67.86, p= ; time: ANOVA F=6.35, p=0.0018). Post-hoc tests did not observe any differences between the two AI conditions in email length and time on task. In contrast, we did observe significant impacts of AI writing style on AI reliance: participants who wrote with the self-assured writing style model relied on the model more (M=0.55, SD=0.34) than participants who wrote with the hesitant writing style model (M=0.45, SD=0.35). These differences were also significant (t=3.16, p=0.0017). We did not observe gender effects or interaction effects between gender and the AI treatment on any of these metrics.
@@ 4.1. Inclusion
We examine the impact of AI writing style on our key measures, beginning with perceptions of inclusion. Figure 2 shows the frequency of different survey responses for each gender condition (different columns) for the inclusion statements: the writing assistant was made for people like me (“Assistant Was Made for You” on the left), and the writing assistant sounded like I would write myself (“Assistant Sounds Like You” on the right). In the condition with the hesitant AI writing style (Figure 2a, left), 65.1% of men (first column) and 63% of WNBSD participants (second column) either agreed or strongly agreed (blue and dark blue bars in the column) that the writing assistant was made for people like them. As the figure also shows, in the self-assured AI writing style group (right half), 68.3% of the men and 68.6% of WNBSD participants either agreed or strongly agreed with this statement. In general, participants more readily agreed with the “made for you” statement (Figure 2a) compared to the “sounds like you” statement (Figure 2b).
In Figure 2b, of the participants who received the hesitant treatment (left half), 42.8% of the men and 45.7% of the WNBSD participants either agreed or strongly agreed that the assistant sounded like how they would write themselves. In the condition with the self-assured AI writing style, on the right half of the figure, 50% of the men and 51.8% of the WNBSD participants either agreed or strongly agreed. A two-way ANOVA analysis revealed no significant effects for gender, AI treatment group, or the interaction between gender and treatment group across both inclusion questions shown in Figure 2.